OpenStudy (anonymous):

PLEASE HELP ME 1. Which is the equation to the line passing through points (2, -7) and (5, -1)? y = 2x - 8 y = - 2x + 9 y = 2x - 11 y = - 2x + 6 2. What is the equation of the line that passes through (12, 4) and is perpendicular to the graph of y = -3 over 4x - 2? y = 4 over 3x - 12 y = 3 over 4x - 5 y = -4 over 3x + 20 y = -3 over 4x + 13

OpenStudy (fibonaccichick666):

So, how do you find the equation of a line?

Nnesha (nnesha):

do you know the slope formula ???

OpenStudy (anonymous):

y = mx + b?

OpenStudy (anonymous):

y=2x-11

Nnesha (nnesha):

ok that is equation what is formula we should use to find slope

OpenStudy (anonymous):

oh, well then how did you get y = 2x - 11?

OpenStudy (anonymous):

(2,-7) y=2(2)-11 y=4-11 y=-7

Nnesha (nnesha):

to find the slope you should use this formula  now plug these value into the formula then solve let me know what you get